Two coins are tossed.Let T the number of tails that occur


a.Construct the probability distribution for the random variable T.

b. Construct The probability histogram.


1
Expert's answer
2021-05-03T04:28:00-0400

Sample space S={HH,HT,TH,TT,}


There are total 4 sample space


The value of T can be 0,1,2,3


"P(T=0)=\dfrac{1}{4}"


"P(T=1)=\dfrac{1}{2}"


"P(T=2)=\dfrac{1}{4}"


(a) Probability Distribution is-




(b) and Histogram is-